OpenStudy (brooke..help00):

Last question. Will fan and medal! @will.h If g(x) = 2(x − 4), find the value of x if g(x) = 20. 32 12 14 10

OpenStudy (jiteshmeghwal9):

g(x)=2(x-4)=20 solve for x

OpenStudy (will.h):

well they said g(x) = 20 and we need to solve for x 20 = 2(x - 4) use distrubutive property 20 = 2x - 8 add 8 to both sides 28 = 2x divide both sides by 2 28/2 = x 14 = x hope that helps

OpenStudy (brooke..help00):

Thank you I got 14:)

OpenStudy (will.h):

good job